OpenStudy (blake57roger):

What is the exact area of a circular patio with a radius of 17 ft?

OpenStudy (blake57roger):

A. 8.5π ft² B. 17π ft² C. 72.25π ft² D. 289π ft²

OpenStudy (isaiah.feynman):



OpenStudy (blake57roger):

@gorv

OpenStudy (johnweldon1993):

Well like isaiah said above...Area of a circle = \(\pi r^2\) So since we know r = 17 we have \[\large Area = \pi 17^2\] \[\large Area = 289 \pi \]

OpenStudy (blake57roger):

Which expression can be used to find the area of a circle with a diameter of 80 cm? A. π · 402 B. π · 160 C. π · 80 D. π · 40